# On the Number of Control Nodes in Boolean Networks with Degree Constraints

### v1

@@ Abstract
This paper studies the minimum control node set problem for Boolean networks (BNs) with degree constraints. The main contribution is to derive the nontrivial lower and upper bounds on the size of the minimum control node set through combinatorial analysis of four types of BNs (i.e., $k$-$k$-XOR-BNs, simple $k$-$k$-AND-BNs, $k$-$k$-AND-BNs with negation and $k$-$k$-NC-BNs, where the $k$-$k$-AND-BN with negation is an extension of the simple $k$-$k$-AND-BN that considers the occurrence of negation and NC means nested canalyzing). More specifically, four bounds for the size of the minimum control node set: general lower bound, best case upper bound, worst case lower bound, and general upper bound are studied, where the general lower bound is a value that is not less than the size of the control node set for any BN, the general upper bound is the maximum value of the size of the minimum control node set for any BN, while the best case upper bound (resp., the worst case lower bound) is the minimum (resp., maximum) value currently found, which is obtained from some BN. By dividing nodes into three disjoint sets, extending the time to reach the target state, and utilizing necessary conditions for controllability, these bounds are obtained, and further meaningful results and phenomena are discovered. Notably, all of the above results involving the AND function also apply to the OR function.
@@ 1 Introduction
On the other hand, many studies focus on finding the minimal set of key variables in the BN and thus controlling them to make the system controllable. Here, controllability [7, 19, 20] means that for any pair of states ${\bf x}^{0},{\bf x}^{T}\in\{0,1\}^{n}$, there exists a control sequence steering the state from ${\bf x}(0)={\bf x}^{0}$ to ${\bf x}(t)={\bf x}^{T}$ for some integer $t$. For example, by using graph theory method, the minimal controllability of conjunctive BN (an important subclass of BN) was explored and a sufficient and necessary condition for controllability of conjunctive BN was derived in [21]. Afterwards, the minimal strong structural controllability problem of BNs was studied in [22], and then an effective method for identifying the minimum set of controlled nodes to achieve global stabilization of large-scale BNs was presented in [23].
However, the above-mentioned studies have not further analyzed the minimum number of control nodes required to achieve controllability of BNs from a theoretical perspective, which is a meaningful and important topic. In this paper, we focus on the minimum control node set problem for BNs with degree constraints, and mainly consider four types of BNs, where the indegree and outdegree of each node are both $k$: (i) $k$-$k$-XOR-BN, where each Boolean function only contains XOR operations, (ii) simple $k$-$k$-AND-BN (resp., simple $k$-$k$-OR-BN), where all Boolean functions are simple AND functions (resp., OR functions) without negations, (iii) $k$-$k$-AND-BN with negation (resp., $k$-$k$-OR-BN with negation), that is an extension of simple $k$-$k$-AND-BN (resp., simple $k$-$k$-OR-BN) considering the occurrence of literal $\overline{x}$, and (iv) $k$-$k$-NC-BN, where each Boolean function is a nested canalyzing function [24, 25].
@@
Generally speaking, the following relationship holds: general lower bound $\leq$ best case upper bound $<$ worst case lower bound $\leq$ general upper bound. If general lower bound $=$ best case upper bound (resp., worst case lower bound $=$ general upper bound), these bounds are called tight or optimal [26]. In this paper, $n$ represents the total number of nodes in a BN, and it is clear that $n$ is a trivial general upper bound. If the worst case lower bound is also $n$, then it is tight.
Table 2: Summary of results, where $n$ represents the total number of nodes in a BN, $t^{*}$ is the minimum time that any initial state ${\bf x}^{0}$ can reach any target state ${\bf x}^{T}$, $\beta(k)=\max\left\{\frac{k-l}{2k-l}n+\frac{1}{2k-l}\left(\sum_{j=1}^{l-2}j\cdot|M_{j}|\right)+\frac{l-1}{2k-l}\left(\sum_{j=l-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right),~{}l=1,2,\ldots,\lfloor\frac{k}{2}\rfloor+1\right\}$, $M_{j}$ is the set of nodes where the literal $x_{i}$ or $\overline{x_{i}}$ appears $j$ times in the $k$-$k$-AND-BN with negation, and for any $k_{1}<k_{2}$, $\sum_{j=k_{2}}^{k_{1}}|M_{j}|=0$.
@@
Theorem 2 2-2-XOR-BNs $\frac{n}{t^{*}}$ $n-t^{*}+1$ General lower bound General upper bound
Remark 2 $k$-$k$-XOR-BNs $\frac{n}{t^{*}}$ $n-t^{*}+1$ General lower bound General upper bound
@@
Theorem 4 $k$-$k$-XOR-BNs ($n=[k(k-1)+1]m$ and $(n\mod k)=0$) $\left[1-\frac{k-1}{k(k^{2}-k+1)}\right]n$ General upper bound
@@
Proposition 7 simple $k$-$k$-AND-BN ($n=(2k-1)m+1$ and $m>1$) $\frac{k-1}{2k-1}(n-1)+1$ Best case upper bound
@@
The main results and contributions of this paper are summarized as follows (see Table 2 for details):
  • By dividing the nodes in a BN into three disjoint sets, a general upper bound on the size of the control node set for $k$-$k$-XOR-BNs is derived, which is $\left[1-\frac{k-1}{k(k^{2}-k+1)}\right]n$. The above result is obtained on the premise that any initial state can reach any target state at $t=2$. Further, we consider reducing the number of control nodes by increasing the time to reach the target state, and for any 2-2-XOR-BN, we find a general upper bound $n-t^{*}+1$ and a general lower bound $\frac{n}{t^{*}}$, where $t^{*}\in[1,n]$ is the minimum time that any initial state ${\bf x}^{0}$ can reach any target state ${\bf x}^{T}$. Moreover, a best case upper bound $k+1$ is also obtained for $k$-$k$-XOR-BNs.
  • Construct a simple $k$-$k$-AND-BN with the minimum size control node set $\frac{k-1}{2k-1}(n-1)+1$, which is the best case upper bound currently. Subsequently, by analyzing the necessary conditions for controllability, a general lower bound $\frac{k-1}{2k-1}n$ for any simple $k$-$k$-AND-BN is derived. It is seen that the worst case lower bound for simple $k$-$k$-AND-BNs is $n$. Furthermore, we focus on the $k$-$k$-AND-BN with negation, which is an extension of simple $k$-$k$-AND-BN considering the occurrence of literal $\overline{x}$, and we note that the general lower bound for any $k$-$k$-AND-BN with negation is $\beta(k)$ (see Table 2). Finally, we find that there is a $k$-$k$-NC-BN with the minimum size control node set $\frac{n}{k}$.
@@ 2 Problem Definition
where every $x_{i}(t)$ takes values in $\{0,1\}$ and $f_{i}:\{0,1\}^{n}\rightarrow\{0,1\}$ is a Boolean function. Here, we use $x_{i}$ to denote the node corresponding to $x_{i}(t)$ and $V=\{x_{1},\ldots,x_{n}\}$ to denote the set of nodes in the BN. Let ${\bf x}(t)=[x_{1}(t),\ldots,x_{n}(t)]$, and use $N(V,F)$ to denote this BN, where $F=\{f_{1},\ldots,f_{n}\}$. For each state vector ${\bf x}$, $({\bf x})_{i}$ denotes the state of the $i$th element of ${\bf x}$. Further, $F({\bf x})$ is used to denote the state vector at the next time step (i.e., ${\bf x}(t+1)=F({\bf x}(t))$), and $G(V,E)$ is used to denote the underlying directed graph, that is, $(x_{j},x_{i})\in E$ if and only if $x_{j}$ is an input node of $x_{i}$.
@@
Given a BN $N(V,F)$, find a minimum cardinality set $U\subseteq V$ such that for any initial state ${\bf x}^{0}$ and any target ${\bf x}^{T}$, there exist control signals $u_{i}(t)$ for all $x_{i}\in U$ that drive $N(V,F)$ from ${\bf x}(0)={\bf x}^{0}$ to ${\bf x}(t)={\bf x}^{T}$ for some $t$.
@@ 3 Boolean networks consisting of XOR functions
Let $N(V,F)$ be a given 2-2-XOR-BN. We assume without loss of generality $V_{1}=\{x_{1},\ldots,x_{m}\}$ (by appropriately permuting indices of $x_{i}$s).
@@
At $t=0$, we let $u_{2}(0)=1$ and let $u_{i}(0)=0$ for the other control nodes. At $t=1$, we let $u_{1}(1)=u_{2}(1)=u_{5}(1)=u_{6}(1)=1$ and $u_{3}(1)=0$. Then, we have the following.
@@
At $t=0$, we let $u_{1}(0)=u_{2}(0)=1$ and let $u_{i}(0)=0$ for the other control nodes. At $t=1$, we let $u_{1}(1)=u_{2}(1)=0$ and $u_{5}(1)=u_{6}(1)=1$. Then, we have the following.
@@
At $t=0$, we let $u_{2}(0)=1$ and let $u_{i}(0)=0$ for the other control nodes. At $t=1$, we let $u_{1}(1)=u_{2}(1)=u_{6}(1)=1$ and $u_{5}(1)=0$. Then, we have the following.
@@
First, for any 2-2-XOR-BN, where $V_{1}^{0}=\emptyset$, we have $V_{1}^{2}=\emptyset$ and $|V_{1}^{1}|=|V_{2}|=m$. If $V_{1}^{0}=\emptyset$, then $|\Gamma^{+}(x_{i})\cap V_{1}|\geq 1$ for any $x_{i}\in V_{1}$. Since $|V_{1}|=m$ and $\Gamma^{+}(x_{i})\cap\Gamma^{+}(x_{j})=\emptyset$ for any $x_{i}\neq x_{j}$ in $V_{1}$, we have $|\Gamma^{+}(x_{i})\cap V_{1}|=1$ for any $x_{i}\in V_{1}$, which means that $V_{1}^{2}=\emptyset$ and $V_{1}=V_{1}^{1}$. And when $|V_{1}|=|V_{1}^{1}|=m$, we have $|V_{2}|=|V_{1}^{1}|=m$ and $|V_{3}|=|V|-|V_{1}^{1}|-|V_{2}|=m$.
@@
$t=1$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ 00 $x_{2}(0)\oplus x_{4}(0)$
$t=1$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ $1$ $x_{2}(0)\oplus x_{4}(0)$
$t=2$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{2}(0)\oplus x_{3}(0)\oplus x_{4}(0)$ 00 $x_{1}(0)\oplus x_{2}(0)$
$t=2$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{2}(0)\oplus x_{3}(0)\oplus x_{4}(0)$ $1$ $x_{1}(0)\oplus x_{2}(0)$
$t=2$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{2}(0)\oplus x_{3}(0)\oplus x_{4}(0)$ 00 $x_{1}(0)\oplus x_{2}(0)$
$t=2$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{2}(0)\oplus x_{3}(0)\oplus x_{4}(0)$ $1$ $x_{1}(0)\oplus x_{2}(0)$
$t=3$ $x_{1}(0)\oplus x_{3}(0)$ $x_{2}(0)\oplus x_{3}(0)$ 00 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $x_{1}(0)\oplus x_{3}(0)$ $x_{2}(0)\oplus x_{3}(0)$ $1$ $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{2}(0)\oplus x_{3}(0)$ 00 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{2}(0)\oplus x_{3}(0)$ $1$ $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{2}(0)\oplus x_{3}(0)}$ 00 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{2}(0)\oplus x_{3}(0)}$ $1$ $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{2}(0)\oplus x_{3}(0)}$ 00 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{2}(0)\oplus x_{3}(0)}$ $1$ $x_{3}(0)\oplus x_{4}(0)$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ 00 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ $1$ $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{4}(0)$ 00 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{4}(0)$ $1$ $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 00 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ $1$ $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 00 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ $1$ $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 00 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ $1$ $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 00 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ $1$ $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ 00 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ $1$ $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{4}(0)$ 00 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{4}(0)$ $1$ $\overline{x_{2}(0)\oplus x_{4}(0)}$
@@
If a 2-2-XOR-BN is controllable, assuming $t^{*}$ is the minimum time for any initial state ${\bf x}^{0}$ to reach any target state ${\bf x}^{T}$, then the following sets are not empty,

### v2

@@ Abstract
This paper studies the minimum control node set problem for Boolean networks (BNs) with degree constraints. The main contribution is to derive the nontrivial lower and upper bounds on the size of the minimum control node set through combinatorial analysis of four types of BNs (i.e., $k$-$k$-XOR-BNs, simple $k$-$k$-AND-BNs, $k$-$k$-AND-BNs with negation and $k$-$k$-NC-BNs, where the $k$-$k$-AND-BN with negation is an extension of the simple $k$-$k$-AND-BN that considers the occurrence of negation and NC means nested canalyzing). More specifically, four bounds for the size of the minimum control node set: general lower bound, best case upper bound, worst case lower bound, and general upper bound are studied. By dividing nodes into three disjoint sets, extending the time to reach the target state, and utilizing necessary conditions for controllability, these bounds are obtained, and further meaningful results and phenomena are discovered. Notably, all of the above results involving the AND function also apply to the OR function.
@@ 1 Introduction
On the other hand, many studies focus on finding the minimal set of key variables in the BN and thus controlling them to make the system controllable. Here, controllability [19, 20, 21, 22, 23, 24] means that for any pair of states ${\bf x}^{0},{\bf x}^{T}\in\{0,1\}^{n}$, there exists a control sequence steering the state from ${\bf x}(0)={\bf x}^{0}$ to ${\bf x}(t)={\bf x}^{T}$ for some integer $t$. For example, by using graph theory method, the minimal controllability of conjunctive BN (an important subclass of BN) was explored and a sufficient and necessary condition for controllability of conjunctive BN was derived in [22]. Afterwards, the minimal strong structural controllability problem of BNs was studied in [25], and then an effective method for identifying the minimum set of controlled nodes to achieve global stabilization of large-scale BNs was presented in [26].
However, the above-mentioned studies have not further analyzed the minimum number of control nodes required to achieve controllability of BNs from a theoretical perspective, which is a meaningful and important topic. In this paper, we focus on the minimum control node set problem for BNs with degree constraints, and mainly consider four types of BNs, where the indegree and outdegree of each node are both $k$: (i) $k$-$k$-XOR-BN, where each Boolean function only contains XOR operations, (ii) simple $k$-$k$-AND-BN (resp., simple $k$-$k$-OR-BN), where all Boolean functions are simple AND functions (resp., OR functions) without negations, (iii) $k$-$k$-AND-BN with negation (resp., $k$-$k$-OR-BN with negation), that is an extension of simple $k$-$k$-AND-BN (resp., simple $k$-$k$-OR-BN) considering the occurrence of literal $\overline{x}$, and (iv) $k$-$k$-NC-BN, where each Boolean function is a nested canalyzing function [27, 28]. Note that considering the degree constraint is reasonable because many biologically relevant Boolean functions have a small number of input variables [29].
@@
Generally speaking, the following relationship holds: general lower bound $\leq$ best case upper bound << worst case lower bound $\leq$ general upper bound. If general lower bound == best case upper bound (resp., worst case lower bound == general upper bound), these bounds are called tight or optimal [30]. In this paper, $n$ represents the total number of nodes in a BN, and it is clear that $n$ is a trivial general upper bound. If the worst case lower bound is also $n$, then it is tight. The main results are summarized in Table 2 and their meanings are discussed in Section 8.
Table 2: Summary of results, where $n$ represents the total number of nodes in a BN, $t^{*}$ is the minimum time that any initial state ${\bf x}^{0}$ can reach any target state ${\bf x}^{T}$, $\beta(k)=\max\left\{\frac{k-l}{2k-l}n+\frac{1}{2k-l}\left(\sum_{j=1}^{l-2}j\cdot|M_{j}|\right)+\frac{l-1}{2k-l}\left(\sum_{j=l-1}^{\lfloor\frac{k}{2}\rfloor}|M_{j}|\right),\;l=1,2,\ldots,\lfloor\frac{k}{2}\rfloor+1\right\}$, $M_{j}$ is the set of nodes where the literal $x_{i}$ or $\overline{x_{i}}$ appears $j$ times in the $k$-$k$-AND-BN with negation, and for any $k_{1}<k_{2}$, $\sum_{j=k_{2}}^{k_{1}}|M_{j}|=0$.
@@
Theorem 2 2-2-XOR-BNs
$\frac{n}{t^{*}}$
$n-t^{*}+1$
General lower bound
General upper bound
Remark 2 $k$-$k$-XOR-BNs
$\frac{n}{t^{*}}$
$n-t^{*}+1$
General lower bound
General upper bound
@@
Theorem 4 $k$-$k$-XOR-BNs (n=[k(k1)+1]mn=[k(k-1)+1]m and $(n\mod k)=0$) $\left[1-\frac{k-1}{k(k^{2}-k+1)}\right]n$ General upper bound
@@
Proposition 7 simple $k$-$k$-AND-BN (n=(2k1)m+1n=(2k-1)m+1 and $m>1$) $\frac{k-1}{2k-1}(n-1)+1$ Best case upper bound
@@ 2 Problem Definition
where every $x_{i}(t)$ takes values in $\{0,1\}$ and $f_{i}:\{0,1\}^{n}\rightarrow\{0,1\}$ is a Boolean function. Here, we use $x_{i}$ to denote the node corresponding to $x_{i}(t)$ and $V=\{x_{1},\ldots,x_{n}\}$ to denote the set of nodes in the BN. Let ${\bf x}(t)=[x_{1}(t),\ldots,x_{n}(t)]$, and use $N(V,F)$ to denote this BN, where $F=\{f_{1},\ldots,f_{n}\}$. For each state vector $\bf x$}, $({\bf x})_{i}$ denotes the state of the $i$th element of $\bf x$}. Further, $F({\bf x})$ is used to denote the state vector at the next time step (i.e., ${\bf x}(t+1)=F({\bf x}(t))$), and $G(V,E)$ is used to denote the underlying directed graph, that is, $(x_{j},x_{i})\in E$ if and only if $x_{j}$ is an input node of $x_{i}$. In addition, if there exists an edge from $x_{j}$ to $x_{i}$, then $x_{j}$ is called an incoming node of $x_{i}$, and $x_{i}$ is called an outgoing node of $x_{j}$. The indegree and outdegree of a node $x_{i}$ are defined as the numbers of edges directed toward $x_{i}$ and from $x_{i}$ to other nodes, respectively.
@@
Given a BN $N(V,F)$, find a minimum cardinality set $U\subseteq V$ such that for any initial state ${\bf x}^{0}$ and any target state ${\bf x}^{T}$, there exist control signals $u_{i}(t)$ for all $x_{i}\in U$ that drive $N(V,F)$ from ${\bf x}(0)={\bf x}^{0}$ to ${\bf x}(t)={\bf x}^{T}$ for some $t$.
@@ 3 Boolean networks consisting of XOR functions
(Proof) Let $N(V,F)$ be a given 2-2-XOR-BN. We assume without loss of generality $V_{1}=\{x_{1},\ldots,x_{m}\}$ (by appropriately permuting indices of $x_{i}$s).
@@
At $t=0$, we let $u_{2}(0)=1$ and $u_{1}(0)=u_{3}(0)=u_{5}(0)=u_{6}(0)=0$. At $t=1$, we let $u_{1}(1)=u_{2}(1)=u_{5}(1)=u_{6}(1)=1$ and $u_{3}(1)=0$. Then, we have the following.
@@
where ${\bf x}^{1}$ and ${\bf x}^{2}$ denote the resulting states at $t=1$ and $t=2$ without any control, ${\bf x}^{3}$ represents the state at $t=1$ when the control signals $u_{2}(0)=1$ and $u_{1}(0)=u_{3}(0)=u_{5}(0)=u_{6}(0)=0$ are applied at $t=0$, ${\bf x}^{4}$ denotes the state at $t=2$ under the same control condition at $t=0$ without applying any control at $t=1$, and ${\bf x}^{5}$ represents the state at $t=2$ when the control signals $u_{2}(0)=1,u_{1}(0)=u_{3}(0)=u_{5}(0)=u_{6}(0)=0$ are applied at $t=0$ and $u_{1}(1)=u_{2}(1)=u_{5}(1)=u_{6}(1)=1,u_{3}(1)=0$ are applied at $t=1$.
@@
At $t=0$, we let $u_{1}(0)=u_{2}(0)=1$ and $u_{5}(0)=u_{6}(0)=0$. At $t=1$, we let $u_{1}(1)=u_{2}(1)=0$ and $u_{5}(1)=u_{6}(1)=1$. Then, we have the following.
@@
where ${\bf x}^{1},\ldots,{\bf x}^{5}$ are defined as in Example 1, except that the control signals at $t=0$ and $t=1$ differ.
@@
At $t=0$, we let $u_{2}(0)=1$ and $u_{1}(1)=u_{5}(1)=u_{6}(1)=0$. At $t=1$, we let $u_{1}(1)=u_{2}(1)=u_{6}(1)=1$ and $u_{5}(1)=0$. Then, we have the following.
@@
where ${\bf x}^{1},\ldots,{\bf x}^{5}$ are defined as in Example 1, except that the control signals at $t=0$ and $t=1$ differ.
@@
(Proof) First, for any 2-2-XOR-BN, where $V_{1}^{0}=\emptyset$, we have $V_{1}^{2}=\emptyset$ and $|V_{1}^{1}|=|V_{2}|=m$. If $V_{1}^{0}=\emptyset$, then $|\Gamma^{+}(x_{i})\cap V_{1}|\geq 1$ for any $x_{i}\in V_{1}$. Since $|V_{1}|=m$ and $\Gamma^{+}(x_{i})\cap\Gamma^{+}(x_{j})=\emptyset$ for any $x_{i}\neq x_{j}$ in $V_{1}$, we have $|\Gamma^{+}(x_{i})\cap V_{1}|=1$ for any $x_{i}\in V_{1}$, which means that $V_{1}^{2}=\emptyset$ and $V_{1}=V_{1}^{1}$. And when $|V_{1}|=|V_{1}^{1}|=m$, we have $|V_{2}|=|V_{1}^{1}|=m$ and $|V_{3}|=|V|-|V_{1}^{1}|-|V_{2}|=m$.
@@
$t=1$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ 0 $x_{2}(0)\oplus x_{4}(0)$
$t=1$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ 11 $x_{2}(0)\oplus x_{4}(0)$
$t=2$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{2}(0)\oplus x_{3}(0)\oplus x_{4}(0)$ 0 $x_{1}(0)\oplus x_{2}(0)$
$t=2$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{2}(0)\oplus x_{3}(0)\oplus x_{4}(0)$ 11 $x_{1}(0)\oplus x_{2}(0)$
$t=2$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{2}(0)\oplus x_{3}(0)\oplus x_{4}(0)$ 0 $x_{1}(0)\oplus x_{2}(0)$
$t=2$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{2}(0)\oplus x_{3}(0)\oplus x_{4}(0)$ 11 $x_{1}(0)\oplus x_{2}(0)$
$t=3$ $x_{1}(0)\oplus x_{3}(0)$ $x_{2}(0)\oplus x_{3}(0)$ 0 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $x_{1}(0)\oplus x_{3}(0)$ $x_{2}(0)\oplus x_{3}(0)$ 11 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{2}(0)\oplus x_{3}(0)$ 0 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{2}(0)\oplus x_{3}(0)$ 11 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{2}(0)\oplus x_{3}(0)}$ 0 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{2}(0)\oplus x_{3}(0)}$ 11 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{2}(0)\oplus x_{3}(0)}$ 0 $x_{3}(0)\oplus x_{4}(0)$
$t=3$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{2}(0)\oplus x_{3}(0)}$ 11 $x_{3}(0)\oplus x_{4}(0)$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ 0 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ 11 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{4}(0)$ 0 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{4}(0)$ 11 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 0 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 11 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 0 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 11 $x_{2}(0)\oplus x_{4}(0)$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 0 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 11 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 0 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $\overline{x_{1}(0)\oplus x_{4}(0)}$ 11 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ 0 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $x_{1}(0)\oplus x_{3}(0)$ $x_{1}(0)\oplus x_{4}(0)$ 11 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{4}(0)$ 0 $\overline{x_{2}(0)\oplus x_{4}(0)}$
$t=4$ $\overline{x_{1}(0)\oplus x_{3}(0)}$ $x_{1}(0)\oplus x_{4}(0)$ 11 $\overline{x_{2}(0)\oplus x_{4}(0)}$
@@
where ${\bf x}^{1},\ldots,{\bf x}^{5}$ are defined as in Example 1, except that the control signals at $t=0$ and $t=1$ differ.
@@
(Proof) If a 2-2-XOR-BN is controllable, assuming $t^{*}$ is the minimum time for any initial state ${\bf x}^{0}$ to reach any target state ${\bf x}^{T}$, then the following sets are not empty,